OpenStudy (anonymous):

Need help with linear approximation Use linear approximation to estimate the following quantities. Choose a value of a to produce a small error. 1/203

OpenStudy (anonymous):

I don't understand this question.

OpenStudy (anonymous):

alright, im here boy

OpenStudy (anonymous):

lemme firgure dis out, hold on.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

You know Calculus?

OpenStudy (anonymous):

10-16 (10 min.) Estimating a cost function. 1. Slope coefficient = =  = = $0.35 per machine-hour Constant = Total cost – (Slope coefficient × Quantity of cost driver) = $5,400 – ($0.35 × 10,000) = $1,900 = $4,000 – ($0.35 × 6,000) = $1,900 The cost function based on the two observations is Maintenance costs = $1,900 + $0.35 × Machine-hours 2. The cost function in requirement 1 is an estimate of how costs behave within the relevant range, not at cost levels outside the relevant range. If there are no months with zero machine-hours represented in the maintenance account, data in that account cannot be used to estimate the fixed costs at the zero machine-hours level. Rather, the constant component of the cost function provides the best available starting point for a straight line that approximates how a cost behaves within the relevant range. i hope this helps!!

OpenStudy (anonymous):

what the hell is this? lol

OpenStudy (anonymous):

or were you looking for more of a straight up answer?? hahah(:

OpenStudy (e.cociuba):

Wat?? Tht makes no sense!

OpenStudy (anonymous):

yes it does

OpenStudy (anonymous):

I was kinda looking for an answer related to my question lol

OpenStudy (e.cociuba):

thts straight off the internet lol :P

OpenStudy (anonymous):

lol that was an example...

OpenStudy (anonymous):

and no i did that myself

OpenStudy (e.cociuba):

I no, but it's a confusing one. Maybe if u made it a bit more legible than he would understand lol

OpenStudy (anonymous):

well lemme try to make this easier... ahahah! hold on!

OpenStudy (anonymous):

can you help me find the linear approximation of 1/203

OpenStudy (e.cociuba):

lol :)

OpenStudy (anonymous):

i get the idea of linear approximation but how do I find an approximation of a constant?

OpenStudy (e.cociuba):

A constant-factor approximation?

OpenStudy (anonymous):

Approximate 1/.203. y = -25(0.203) + 10 = -5.075 + 10 = -4.925

OpenStudy (anonymous):

is that correct i think...

OpenStudy (anonymous):

idk where did you get all these other numbers?

OpenStudy (anonymous):

lol... you gotta do the work..

OpenStudy (anonymous):

i give up :(

OpenStudy (anonymous):

sorry, wished i couldve helped!

OpenStudy (anonymous):

ummmmmmmmm okay thanks anyway

OpenStudy (anonymous):

i got u one sec

OpenStudy (anonymous):

ok thnks

OpenStudy (anonymous):

You wish to find a "good" linear approximation of \(f(x)=\frac1x\) near \(x=203\) in order to approximate the value at \(x=203\) using relatively simple computation. How about we decide to pick our approximation at \(a=200\)? \[y-f(a)=f'(a)[x-a]\\y-\frac1{200}=-\frac1{200^2}[x-200]\] Thus we may approximate \(f(203)\approx y(203)=-\frac3{200^2}+\frac1{200}=\frac{200-3}{200^2}=\dots\)

OpenStudy (anonymous):

ok I see what they're doing. Thanks, now it makes senses.